Your firm received an RFP (request for proposal) on a wire harness from GM with fixed costs of $1 million and MC of $1 with expected sales of 1 million units. What is break-even price?
Answer #1

Answer

Break-even occurs when Total revenue(TR) = Total Cost(TC)

Total cost(TC) = Total Fixed cost(TFC) + Total Variable cost(TVC)

Here TFC = 1 million. Also as marginal cost is constant thus TVC = MC*Q.

It is given that MC = marginal cost - 1 and Q = quantity or sales = 1 million => TVC = 1*1 million = 1 million

Thus TC = TVC + TFC

=> TC = 1 million + 1 million = 2 million

TR = Total revenue = P*Q where P = Price and Q= quantity or sales = 1 million

=> TR = (1 million)*Q

Thus TR = TC => (1 million)*P = 2 million

=> P = $2

Hence Break-even price is $2